OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation 1 Equation 2 6-x=5x+30 4x-8=3x+2 A. The solution to equation #1 is smaller than the solution to equation #2. B. The solution to equation #1 is larger than the solution to equation #2. C. The solution to equation #1 is the same as the solution to equation #2. D. None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

@Jacob902

Vocaloid (vocaloid):

start with equation 1 6-x = 5x+30 add x to both sides, what do you get?

OpenStudy (jacob902):

Eqn (1) 6 - x = 5x + 30 ==> 4x = - 24 ==> x = - 6 Eqn (2) 4x - 8 = 3x + 2 ==> x = 10 The solution to equation #1 is smaller than the solution to equation #2. is true

OpenStudy (anonymous):

1x-x? idk

OpenStudy (anonymous):

so A

OpenStudy (jacob902):

correct

OpenStudy (anonymous):

thnx :D